# Selection rule of canonical differential equations from Intersection theory

### v1

@@ 3.3 Condition of $\big{(}\eta^{-1}\big{)}_{JK}\neq 0$
Notice that $\eta_{JK}$ is non-zero only when $\varphi_{J;\hat{K}}$ and $\varphi_{K;\hat{K}}$ share $n_{\hat{J}}$-SP and each of them gives contribution $C_{J}^{(\bm{-1})}C_{K}^{(\bm{-1})}/\bm{\gamma}^{(\alpha)}_{\hat{K}}$. Also notice that $\big{(}\eta^{-1}\big{)}_{JK}=(-1)^{J+K}|\eta^{(KJ)}|/|\eta|$, in which $\eta^{(KJ)}$ is the adjugate matrix of $\eta$. Then follow the similar analysis as [40], we need modify the conception of "n-SP chain" in [40], which is unoriented, to the following one
  • If $\dot{\varphi}_{I;\hat{J}}$ and $\varphi_{J;\hat{J}}$ share $n_{\hat{J}}$-SP, we say $\varphi_{I}$ is Maximal multivariate Simple Pole (MSP) related to $\varphi_{J}$, and denoted as $\varphi_{I}\rightarrow\varphi_{J}$ or $\varphi_{J}\leftarrow\varphi_{I}$. Now, it is an orient relation. $\varphi_{I}\rightarrow\varphi_{J}$ does not imply $\varphi_{J}\rightarrow\varphi_{I}$. "Maximal" for $\dot{\varphi}_{I;\hat{J}}$ and $\varphi_{J;\hat{J}}$ are $n_{\hat{J}}$-forms and share $n_{\hat{J}}$-SP.
  • If $\varphi_{I}\rightarrow\varphi_{J}$, we also say that $\varphi_{I}$ is linked to $\varphi_{J}$ via a MSP chain. If $\varphi_{I}\rightarrow\varphi_{J}$ and $\varphi_{J}\rightarrow\varphi_{K}$, then we say $\varphi_{I}$ is linked to $\varphi_{K}$ via the MSP chain $\varphi_{I}\rightarrow\varphi_{J}\rightarrow\varphi_{K}$, we denote it as $\varphi_{I}\rightarrow\rightarrow\varphi_{J}$. Similar for more forms.
Then, only when $\varphi_{J}\rightarrow\rightarrow\varphi_{K}$, $\big{(}\eta^{-1}\big{)}_{JK}$ could be non-zero.
@@ 3.4 Conclusion
Only when there exist at least a $\varphi_{J}$, satisfy that $\varphi_{J;\hat{J}}$ share at least one $n_{\hat{J}}$-SP or ($n_{\hat{J}}$-1)-SP with $\varphi_{I;\hat{J}}$ and $\varphi_{J}\rightarrow\rightarrow\varphi_{K}$, then $\Omega_{IK}$ could be non-zero. When all $\beta_{i}$ in $u=\prod_{i}\left[P_{i}(\bm{z})\right]^{\beta_{i}}$ are proportional to $\epsilon$, $\Omega_{IK}$ is canonical form.
4 Univariate Example: Compare Regulator with Relative Cohomology
Consider $u=z^{\delta}(z-c_{1})^{\epsilon}(z-c_{2})^{\epsilon}$ with regular $\delta$. Constructing the $\mathrm{d}\log$ basis as
@@ 4.1 Regulator method
If compute the differential equations of the basis via regulator method, we have
@@
Then, $\eta_{11}$ receive contribution from shared 1-SP at $p_{1}=0$ and $p_{4}=\infty$ of $\varphi_{1}$ itself, so it equals $1/\gamma_{1}+1/\gamma_{4}$. Similar analysis also could be applied to other $\eta_{JK}$ and give
@@
Let’s calculate $\braket{\dot{\varphi}_{1}}{\varphi_{1}}$ as example. For the contribution from the shared 1-SP $p_{1}=0$, one could use (15) and get
@@
where $t=1/z$. Thus it does not contribute. Continue the progress, we will have
@@
It is redundant and $\hat{\mathrm{d}}\Omega^{\text{reg}}_{21}$ vanish when $\delta=0$, but the selection rules get from the regulator method in [40] could not tell you that.
@@ 4.2 Relative cohomology method
If we calculate the $\Omega$ via relative cohomology method, we have
@@
Calculation of $\braket{\varphi_{I}}{\Delta_{\hat{1}}\varphi_{1}}$ will be simplified, become we need apply the cut of $\mathcal{B}_{\hat{1}}$:
@@
We have $\hat{\mathrm{d}}\Omega=\hat{\mathrm{d}}\Omega^{\text{reg}}(c_{i};\epsilon,0)$. It shows that the relative cohomology method simplifies the computation, avoids redundancy of regulators, and gets the exact selection rules of CDE as it could tell you $\Omega_{21}=0$ here.
@@ 5 Summary and Outlook
We improve the selection rules of CDE, provide the $\mathrm{d}\log$ projection perspective, and avoid redundancy via relative cohomology. The mathematical structure we have found may benefit many aspects in evaluating perturbation quantum field theory, such as reduction (since CDE themselves are reduction relations), symbology bootstrap, and analytic solution of Feynman integrals.

### v2

@@ 3.3 Condition of $\big{(}\eta^{-1}\big{)}_{JK}\neq 0$
Notice that $\eta_{JK}$ is non-zero only when $\varphi_{J;\hat{K}}$ and $\varphi_{K;\hat{K}}$ share at least an $n_{\hat{J}}$-SP and each shared $n_{\hat{J}}$-SP contributes $C_{J}^{(\bm{-1})}C_{K}^{(\bm{-1})}/\bm{\gamma}^{(\alpha)}_{\hat{K}}$. Also, note that $\big{(}\eta^{-1}\big{)}_{JK}=(-1)^{J+K}|\eta^{(KJ)}|/|\eta|$, where $\eta^{(KJ)}$ is the adjoint matrix of $\eta$. Following a similar analysis as in [40], we need to modify the concept of "n-SP chain" from [40], which is unoriented, to a new concept of cut-$n$-SP chain:
  • If ${\varphi}_{I;\hat{J}}$ and $\varphi_{J;\hat{J}}$ share $n_{\hat{J}}$-SP, we say $\varphi_{I}$ is cut-$n$-SP related to $\varphi_{J}$, and denoted as $\varphi_{I}\rightarrow\varphi_{J}$ or $\varphi_{J}\leftarrow\varphi_{I}$. Now, it is an orient relation, i.e., $\varphi_{I}\rightarrow\varphi_{J}$ does not imply $\varphi_{J}\rightarrow\varphi_{I}$.
  • If $\varphi_{I}\rightarrow\varphi_{J}$, we also say that $\varphi_{I}$ is linked to $\varphi_{J}$ via a cut-$n$-SP chain. If $\varphi_{I}\rightarrow\varphi_{J}$ and $\varphi_{J}\rightarrow\varphi_{K}$, then we say $\varphi_{I}$ is linked to $\varphi_{K}$ via the cut-$n$-SP chain $\varphi_{I}\rightarrow\varphi_{J}\rightarrow\varphi_{K}$, we denote it as $\varphi_{I}\rightarrow\rightarrow\varphi_{J}$. Similar understanding for more forms $\varphi$.
Then, $\big{(}\eta^{-1}\big{)}_{JK}$ could be non-zero only when $\varphi_{J}\rightarrow\rightarrow\varphi_{K}$.
@@ 3.4 Conclusion
With the above discussions, we have CDE selection rules: If all master integrals have $\mathrm{d}\log$-form $\varphi_{I}$s, the differential equations $\big{(}\hat{\mathrm{d}}\Omega\big{)}_{IK}$ is $\mathrm{d}\log$-form and given by the simple formula of LO contribution of intersection number. When all $\beta_{i}$ in eq.(2) are proportional to $\epsilon$, $\big{(}\hat{\mathrm{d}}\Omega\big{)}_{IK}$ is reduced to be proportional to $\epsilon$, thus is canonical form. $\big{(}\hat{\mathrm{d}}\Omega\big{)}_{IK}$ could be non-zero only when there exist a $\varphi_{J}$ that $\dot{\varphi}_{I;\hat{J}}$ and $\varphi_{J;\hat{J}}$ share $n_{\hat{J}}$-SP or ($n_{\hat{J}}-1$)-SP, and $\varphi_{J}\rightarrow\rightarrow\varphi_{K}$.
4 Univariate Example: Comparing Regulator Method with Relative Cohomology
Consider $u=z^{\delta}(z-c_{1})^{\epsilon}(z-c_{2})^{\epsilon}$, where $\delta$ is regulator. We construct the $\mathrm{d}\log$ basis as follow:
@@ 4.1 Regulator method
If we compute the differential equations of the basis using the regulator method, we obtain:
@@
Then, $\eta_{11}$ receives contributions from shared 1-SP at $p_{1}=0$ and $p_{4}=\infty$ of $\varphi_{1}$ itself, so it equals $1/\gamma_{1}+1/\gamma_{4}$. A similar analysis can be applied to other $\eta_{JK}$, giving:
@@
Let’s calculate $\braket{\dot{\varphi}_{1}}{\varphi_{1}}$ as an example. For the contribution from the shared 1-SP $p_{1}=0$, we use (15) to get
@@
where $t=1/z$. Thus, this term does not contribute. Continuing the calculation, we find:
@@
It is redundant and $\hat{\mathrm{d}}\Omega^{\text{reg}}_{21}$ vanishes when $\delta=0$, but the selection rules get from the regulator method in [40] do not indicate this.
@@ 4.2 Relative cohomology method
If we calculate $\Omega$ using relative cohomology method, we have
@@
Calculation of $\braket{\varphi_{I}}{\Delta_{\hat{1}}\varphi_{1}}$ is simplified as we apply the cut of $\mathcal{B}_{\hat{1}}$:
@@
We have $\hat{\mathrm{d}}\Omega=\hat{\mathrm{d}}\Omega^{\text{reg}}(c_{i};\epsilon,0)$. This demonstrates that the relative cohomology method simplifies the computation, avoids redundancy from regulators, and provides a more precise version of the selection rules for CDE. In particular, it correctly shows that $\Omega_{21}=0$ in this case.
@@ 5 Summary and Outlook
We improve the selection rules for CDE by providing a perspective from the $\mathrm{d}\log$ projection and avoiding redundancy through the relative cohomology method. The mathematical structure we have developed can benefit various aspects of perturbative quantum field theory, including reduction (since CDEs are themselves reduction relations), symbolic bootstrap, and the analytic evaluation of Feynman integrals.